Each month, a Filipino household generates an average of 28 pounds of newspaper for garbage or recycling. Assume the standard deviation is 2 pounds. If a household is selected at random, find the probability of its generating more than 31 pounds. (

1
Expert's answer
2022-03-20T13:24:48-0400

"P(X>31)=P(Z>\frac{31-28}{2})=P(Z>1.5)=1-P(Z<1.5)=0.0688."